2 point charges are placed on x axis.  a 2μc charge at x=10cm and -1μc at x=40cm. Calculate the potential at point x=100cm
 

Grade:12

1 Answers

venkat
6 years ago
First you have to find the potential due to first charge and potential due to second charge at x=100cm and add them algebrically.
V_{1}=q/4\pi \epsilon _{0}r
V_{1}=2*10^{-6}*9*10^{9}/90*10^{-2}
V_{1}=18*10^{3}/9*10^{-1}
V_{1}=2*10^{4}V
Similarly find out V2
You will get
V_{2}=-1.5*10^{4}V
Net potential at x=100cm is
V=V1+V2
V=0.5*104V
V=5*103V